# Dynamic planning in hierarchical active inference

### v2

@@ 5 Discussion
Despite the many advances that have been made in this relatively new and promising research area, with increasing popularity among different scientific domains, a current drawback is that studies about low-level motor control and high-level behavior have been somewhat distinct so far, making use of two highly specular but separated frameworks. As a result, there is no consensus on how to achieve dynamic planning (i.e., how to perform decision-making in constantly changing environments), and state-of-the-art solutions to tackle complex tasks generally couple active inference with traditional machine learning methods. From a theoretical perspective, a few works prescribed an efficient and elegant way for combining the capabilities of discrete and continuous representations into a single generative model [48, 59]; however, this hybrid approach has not reached as much maturity, with the consequence that there are far fewer studies on the subject in the literature, none applied to dynamic contexts.
@@
For this reason, we tried here to give a comprehensive view of this yet unexplored direction, comparing several design choices regarding goal-directed behavior, with the intent of bringing motor control and behavioral studies closer. As a practical goal, we decided to model tool use [99], a task that inevitably calls for both discrete and continuous frameworks, and that requires taking two additional aspects into account, i.e., object affordances and hierarchical causal relationships. In a simple scenario, considering a target to reach as the cause of some hidden states is a reasonable assumption and makes the agent able to operate in dynamic contexts. But assuming that multiple objects are present, how does the agent decide which one will be the cause of a particular action? And what if the target moves along a non-trivial path? If the hidden states are factorized into independent distributions encoding multiple entities, the hidden causes may be seen from a different angle, i.e. they would manipulate the hidden states through flexible intentions [43, 44, 75]. Each of these entities would have its own dynamics, allowing the agent to predict, e.g., the trajectory of a moving ball. Then, this unit was scaled up to construct complex (deep) hierarchical structures, e.g., for simulating human body kinematics [44], and to perform more general transformations of reference frames, e.g., perspective projections [58]. A hierarchical factorization of the hidden states now assumes a broader perspective that can also account for multi-agent interaction – an aspect that has been analyzed in the discrete framework as well [116]. Finally, we designed a hybrid unit with discrete hidden causes and continuous hidden states, affording dynamic inference via Bayesian model reduction [69] that, when coupled with a higher-level discrete model, made it possible to simulate multi-step tasks involving online planning of actions. This showed further parallelisms between the inference of intentions in the continuous domain, and policies in discrete models.
@@
Still, the real question is how we can reach such performances without embedding our prior knowledge into the agent’s generative model. Although not relevant for many continuous-time implementations that focused on different aspects of motor control, a common criticism is that the structure of generative models is a-priori defined and fixed, with intricate and hardcoded dynamics functions that raise some concerns about biological plausibility. In contrast, one appealing characteristic of PCNs is that they simulate brain processing with extremely simple functions typical of the connectivity of neural networks (e.g., linear combinations of weights and biases passed to a non-linear activation function). This allows the network to easily adapt to high-dimensional data, with a few critical advantages compared to deep learning arising from a top-down modulation [16]. While much of this research involves static representations, some studies began to address how predictive coding could be used to learn temporal sequences [117, 118], or to solve RL tasks [119, 103, 120]. Here, we demonstrated how generative models in active inference could be realized by simple likelihood and dynamics functions, showing some analogies with the inferential process of PCNs. Based on these findings, a promising research direction would be to imitate their multiple-input and multiple-output architectures (as in Figure 9(d)), so that an agent could not only learn – in a biologically plausible way – its kinematic configuration and the system dynamics, but also act over them to conform to prior beliefs.
@@
Learning policies in continuous environments is not an easy challenge, but addressing it with strategies different from traditional methods might be key for advancing with current intelligent agents, realizing the full theoretical potential at the basis of active inference and the free energy principle. On this matter, the state-of-the-art is to approximate the likelihood and transition distributions by deep neural networks [78, 121, 122, 123]. While several benefits arise compared to deep RL, this still relegates the deep structure within the neural network, generally making use of a single-level active inference agent. One study used a more biologically plausible PCN as a generative model [120], but relied on a similar approach. As extensively analyzed in [51], neural networks can be seen as static generative models with infinitely precise priors at the last level and no hidden states. This architecture can be used to perform sparse coding or Principal Components Analysis (PCA); however, it fails to account for dynamic variables, as in deconvolution problems or filtering in state-space models. Temporal depth – either discrete or continuous – is thus key to inferring the most accurate representation of the environment, and indeed it seems that cortical columns are able to express model dynamics (e.g., the prefrontal cortex is constantly involved in predicting future states, and motion-sensitive neurons have been recorded in the early visual cortex as well [124]). While it is true that temporal sequences can be easily handled by deep architectures such as recurrent neural networks or transformers [125], their passive generative mechanism could still reflect to the behavior of the active inference agent. In contrast to such a passive AI, being grounded on sensorimotor experiences and actively modifying the environment could be fundamental to the emergence of genuine understanding [126]. Taken together, these facts suggest that acting upon generalized coordinates of motion or discrete future states also for intermediate levels could bring several advantages in solving RL tasks. For instance, representing an agent in a hierarchical fashion afforded highly advanced control over its whole body structure that would not have been possible by a single level generating only the hand position [44, 115].
@@
How then to learn dynamic planning in deep hierarchical models? In [127], it has been stressed the importance of being discrete when considering structure learning. Indeed, hierarchical discrete models afford much more expressivity compared to their continuous counterparts, above all, deriving from the simplicity of computing the expected free energy that allows an agent to plan actions over the imminent future. Nonetheless, as Friston and colleagues note, whether using continuous or discrete representations depends on the model evidence. Specifically, the former may have better performances when the evidence has contiguity properties, e.g., when dealing with time series or with Euclidean space. Indeed, the task exemplified in Figure 18 is effective because the Bayesian model reduction performs a dynamic evidence accumulation over the extrinsic space in which the agent operates. Hence, coupling the hierarchical depth of the hybrid units in Figure 17 with a hierarchical discrete architecture (and not just a single discrete level) could bring efficient structure learning also in constantly changing environments. An alternative approach would be to combine in a hierarchical fashion units composed of a joint discrete-continuous model – as in Figure 15 – which would allow to perform dynamic planning within each single unit. While this solution may not be supported yet by empirical evidence from biological agents, it could be an encouraging direction to explore from a machine learning perspective, contrasting the hypothesis of central discrete decision-making with a distributed network of local decisions.
@@
A third interesting topic regards motor intentionality. Although multi-step tasks are typically tackled at the discrete level, we showed here that, under appropriate assumptions, a non-trivial behavior could be achieved and analyzed also at the continuous level. The flexible intentions that we defined could be compared to an advanced stage of motor skill learning, consisting of autonomous and smooth movements that do not necessitate conscious decision-making [75]. Still, the model structure was predefined in this case as well. How do such intentions emerge during repeated exposure to the same task? How does the agent score which intentions will be appropriate for a specific context? As mentioned in the last chapter, the optimization of intention precisions is likely to involve the free energy of reduced models (see Equation 38). This process may shed light on how discrete actions arise from low-level continuous intentions and, conversely, how the latter are generated from a composite discrete action. Last, a few studies proposed additional connections between policies unfolding at different timescales, either directly [111, 112] or through discrete hidden states [110]. Such approaches could be adopted in hybrid and continuous contexts as well, so that flexible intentions could be propagated via local message passing between hidden causes along the whole hierarchy.
6 Acknowledgments

### v3

@@ 5 Discussion
Despite the many advances that have been made in active inference, with increasing popularity among different scientific domains, a current drawback is that studies about motor control and decision-making have been somewhat separated so far, making use of two highly similar but distinct frameworks. As a result, there is no consensus on how to achieve dynamic planning (i.e., how to perform decision-making in constantly changing environments), and state-of-the-art solutions to tackle complex tasks generally couple active inference with traditional methods in machine learning or optimal control. From a theoretical perspective, a few works prescribed an efficient and elegant way for combining the capabilities of discrete and continuous representations into a single generative model [49, 67]; however, this hybrid approach has not reached as much maturity, with the consequence that there are far fewer studies on the subject in the literature.
@@
For this reason, we tried here to give a comprehensive view of this yet unexplored direction, comparing several design choices regarding tasks of increasing complexity, with the intent of bringing motor control and behavioral studies closer. As a practical example, we described the modeling of tool use [90], a task that inevitably calls for both discrete and continuous frameworks, and that requires taking two additional aspects into account, i.e., object affordances and hierarchical relationships. In a simple scenario, considering a target to reach as the cause of some hidden states is a reasonable assumption and permits operating in dynamic contexts. But when there are multiple objects, how does the agent decide what the cause of a particular action will be? How can it account for different object affordances? And what if the target moves along a non-trivial path? The hidden states may be factorized into independent distributions encoding multiple entities in intrinsic coordinates, hence expressing potential body configurations. Further, the hidden causes may be linked to potential trajectories related to the agent’s intentions [64, 80, 66]. Beliefs over each entity would have their own dynamics, allowing the agent to predict, e.g., the trajectory of a moving ball. Then, we described how this unit could be scaled up to construct complex (deep) hierarchical structures, e.g., for simulating human body kinematics [80], and to perform more general transformations of reference frames, e.g., perspective projections [83]. A hierarchical factorization of the hidden states now assumes a broader perspective that can also account for multi-agent interaction – an aspect that has been analyzed in the discrete framework as well [130]. Finally, we described the design of a hybrid unit with discrete hidden causes and continuous hidden states, affording dynamic inference via Bayesian model reduction [65]; then, a higher-level discrete model made it possible to simulate dynamic tasks involving online planning of actions. This showed further parallelisms between the inference of trajectories in continuous models and policies in discrete models.
One challenge in maintaining a deep representation of the kinematic chain is the associated computational complexity and the time required to infer a body posture from visual input, both increasing with the DoF. This is because an extrinsic prediction error generated at the distal (e.g., hand) levels would have to climb the hierarchy toward the (root) body-centered reference frame. This directly affects behavioral accuracy and movement time, which critically depend on the correct inference of the intrinsic state. A performance comparison during inference and action is shown in [80], for increasing DoF. In contrast, handling multiple objects does not lead to a significant increase in inference time, since the (deep) representations are computed in parallel, and are only limited by the complexity of the object dynamics. Communication with the discrete model is equally efficient, as a single discrete state can infer the body trajectories based on multiple units concurrently [65]. Still, the behavior of other objects in the environment requires predicting not only how the agent will behave after a specific action but also how the other objects will behave and their influence on the agent’s behavior. This requires a rich discrete representation that can model the interactions of every object, and evaluating complex policies may not be as effective for high-dimensional settings [131]. A last issue regards the time $T$ needed for accumulating continuous evidence for a single discrete step $\tau$. As shown in [66], Figure 5b, a narrow sampling time interval permits efficient control in highly dynamic environments but at the cost of increased computational time.
@@
A limitation of the models reviewed here is their fixed structure. The critical question, then, is how tool use can be realized without embedding prior knowledge into the agent’s generative model. In other words, how an agent can perform using active inference when starting from a blank memory or under the assumption that the surrounding environment will remain unchanged. A common criticism of continuous-time active inference models is that their generative models are a-priori defined and fixed, with intricate and hardcoded dynamics functions that raise some concerns about biological plausibility. In contrast, one appealing characteristic of PCNs is that they simulate brain processing with extremely simple functions typical of the connectivity of neural networks (e.g., linear combinations of weights and biases passed to a non-linear activation function). This allows PCNs to easily adapt to high-dimensional data, with a few critical advantages compared to deep learning (e.g., about top-down modulation) [25]. While much of the research with PCNs involves static representations, some studies addressed how predictive coding could be used to learn temporal sequences [132, 89], or to solve RL tasks [133, 91, 134]. Here, we demonstrated how generative models in active inference could be realized by simple likelihood and dynamics functions, showing some analogies with the inference of PCNs. Based on these findings, a promising research direction would be to imitate their (deep) hierarchical architectures (as in Figure 9(d)), so that an agent could not only flexibly adapt its body schema to interact with objects with different hierarchical structures, but also learn the system dynamics and act over them to conform to prior beliefs.
@@
Learning policies in continuous environments is not an easy challenge, but addressing it with strategies different from traditional methods might be key for advancing with current intelligent agents, realizing the full theoretical potential at the basis of active inference and the free energy principle. On this matter, the state-of-the-art is to approximate the likelihood and transition distributions by deep neural networks [135, 136, 137, 138, 139, 140, 141, 142]. While several benefits arise compared to deep RL, this still relegates the deep structure within the neural network, generally using a single-level active inference agent. One study used a more biologically plausible PCN as a generative model [134], but relied on a similar approach. As extensively analyzed in [57], neural networks can be seen as static generative models with infinitely precise priors at the last level and no hidden states. This architecture can be used to perform sparse coding or Principal Components Analysis (PCA); however, it fails to account for dynamic variables, as in deconvolution problems or filtering in state-space models. Temporal depth – either discrete or continuous – is thus key to inferring the most accurate representation of the environment, and indeed it seems that cortical columns are able to express model dynamics (e.g., the prefrontal cortex is constantly involved in predicting future states, and motion-sensitive neurons have been recorded in the early visual cortex as well [143]). While it is true that temporal sequences can be easily handled by deep architectures such as recurrent neural networks or transformers [144], their passive generative mechanism could still be reflected in the behavior of the active inference agent. In contrast to such a passive AI, being grounded on sensorimotor experiences and actively modifying the environment could be fundamental to the emergence of genuine understanding [145]. Taken together, these facts suggest that a hierarchy of actions upon generalized coordinates of motion or discrete future states could bring several advantages in solving RL tasks. For instance, representing an agent in a hierarchical fashion afforded highly advanced control over its whole body structure that would not have been possible by a single level generating only the hand position [80, 129].
@@
How to learn dynamic planning in deep hierarchical models? The importance of being discrete when considering structure learning has been stressed in [146]. Indeed, hierarchical discrete models afford much more expressivity compared to their continuous counterparts, above all, deriving from the simplicity of computing the expected free energy. Nonetheless, as Friston and colleagues note, whether using continuous or discrete representations depends on the model evidence. Specifically, the former may have better performances when the evidence has contiguity properties, e.g., when dealing with time series or with Euclidean space. Indeed, the task exemplified in Figure 18 is effective because the Bayesian model reduction performs a dynamic evidence accumulation over the extrinsic space in which the agent operates. Hence, coupling the hierarchical depth of the hybrid units in Figure 17 with a hierarchical discrete architecture (and not just a single discrete level) could bring efficient structure learning also in constantly changing environments. A successful Bayesian approach to learning discrete structures involves growing discrete distributions using an infinite Dirichlet process and the Chinese restaurant prior [147]. This method assumes a potentially infinite mixture of basis distributions and builds a structure starting from an empty model wherein novel configurations are either assigned to popular existing states, or occasionally, used to create novel states. This approach has shown success in learning structures that support complex goal-directed behavior [148], hierarchical spatial organization [17], and spatial navigation [16] among other applications. An alternative to hierarchical discrete models would be to combine units composed of a joint discrete-continuous model – as in Figure 15 – which would allow to perform dynamic planning within each single unit. While this solution may not be supported by empirical evidence from biological agents, it could be an encouraging direction to explore from a machine learning perspective, contrasting the hypothesis of central discrete decision-making with a distributed network of local decisions.
@@
A third interesting topic regards motor intentionality. Although multi-step tasks are typically tackled at the discrete level, we showed here that, under appropriate assumptions, a non-trivial behavior could be achieved and analyzed also at the continuous level. The flexible intentions we defined could be compared to an advanced stage of motor skill learning, consisting of autonomous and smooth movements that do not necessitate conscious decision-making [66]. Still, even in this case the model structure was predefined. How do such intentions emerge during repeated exposure to the same task? How does the agent score which intentions will be appropriate for a specific context? As mentioned in the last chapter, optimization of dynamics precisions is likely to involve the free energy of reduced models (see Equation 38). This process may shed light on how discrete actions arise from low-level continuous trajectories and, conversely, how the latter are generated from a composite discrete action. Last, a few studies proposed additional connections between policies unfolding at different timescales, either directly [124, 125] or through discrete hidden states [123]. Such approaches could be adopted in hybrid and continuous contexts as well, so that flexible intentions could be propagated via local message passing between hidden causes along the whole hierarchy.
Data availability
Code and data are deposited in GitHub (https://github.com/priorelli/dynamic-planning).
Acknowledgments